# Watermarking Language Models through Language Models

### v1

@@ 5 Conclusion
In this paper, we proposed a dynamic framework for watermarking LMs using a multi-model setup that includes a Prompting LM, a Marking LM, and a Detecting LM. Our primary objective was to create an adaptive watermarking process that leverages model-generated prompts to embed and detect watermarks across diverse LLMs, offering a scalable solution for content attribution and model authenticity verification.
Through a series of experiments, we evaluated the effectiveness of our framework with both ChatGPT and Mistral models. Initial experiments demonstrated the framework’s capacity to embed and detect watermarks with high accuracy, achieving 95% accuracy for ChatGPT-generated responses and 88.79% for those from Mistral. A combined dataset experiment further assessed whether a unified Detecting LM could generalize across different Marking LMs; however, results indicate that model-specific training may yield better performance, aligning with our original design assumptions.

### v2

@@ 5 Conclusion
In this work, we introduced a novel prompt-driven framework for watermarking large language models that leverages natural instructions to embed imperceptible but detectable signals in generated outputs. Unlike prior methods that rely on fixed token manipulations or access to model internals, our approach uses a multi-model design—comprising a Prompting LM, a Marking LM, and a Detecting LM—to support flexible, black-box-compatible watermarking across diverse architectures.
Extensive experiments across over 25 Prompting–Marking LM combinations demonstrate that our framework consistently achieves high watermark detection accuracy, even under realistic conditions such as fine-tuning, distillation, and adversarial prompt-based attacks. These results underscore the robustness, modularity, and generalizability of prompt-based watermarking. Moreover, our ablation analysis reveals how different linguistic strategies—semantic, lexical, and structural—affect watermark signal strength and detectability, providing insights into the design of effective watermarking instructions.
By treating prompts as instruction-bearing control channels, our framework offers a scalable and adaptive approach to content attribution, model authentication, and responsible LLM deployment. Future work can explore joint optimization of watermark generation and detection, expand to multilingual settings, and integrate cryptographic properties for stronger security guarantees.